Attendees
Attendees | Regrets | Notified |
---|---|---|
Agenda
News
- Power outage at DSOC/NMTech Tuesday pm due to windstorm, mostly recovered, but many VLASS-related tests delayed. Nevertheless D-config processing has started.
- Program Administration
- Science Ready Data Products
- AAT4.3.4/WS2.8.7/VLASS 2.4.4 deployed
- Enables cal+img, easy to change default workflow processing settings.
- Noted that is only running serial (4-way MPI was requested), not a problem in D-config (in fact may be better in terms of efficiency as cal can only run serial), but may need to fix for A & B configs to prevent overly-long runs.
- Mitigation may be adequate if implemented, leave at serial for now (more efficient overall).
- Enables cal+img, easy to change default workflow processing settings.
- Pipeline CASA 6.6.1-17-pipeline-2024.1.1.22 now installed
- Some hiccups in the cross-site deployment of CASA would have delayed us if power outage hadn't happened... Need improved coordination between SIS and CASA, better communication with NMOps.
- Discuss with Laura Jensen how this might be improved in future.
- default recipe for AUDI will try to reapply existing selfcal rather than redo, could handle these as re-runs if needed or do a patch.
- Decided to rerun pending refactoring of AUDI later this year.
- QA-related pipeline requests in the "no flagging" era
- Large image/cube mitigation. PIPE-2452 - Getting issue details... STATUS
- AQUA report PIPE-730 - Getting issue details... STATUS
- Tickets to be created: VLA QA Evaluation criteria
- Suggest "official" handover of SRDP VLA Ops March 31st. From then, Amy will be responsible for tracking statistics, VLA sections of SRDP webpages, SRDP criteria for VLA data.
- Maybe write an e-News article - spectral line pipeline, 3C48 back as an SRDP calibrator, transfer to NMOps etc.
- For next month's enews and focus on user-facing changes.
- OK to go from "SRDP" to "Quality assessed"
- AAT4.3.4/WS2.8.7/VLASS 2.4.4 deployed
- VLASS
- Planning for VLASS-II
- Now have Texas TACC account set up.
- Plan sent to Jeff for DMS effort request for VLASS-II, no response yet
- JHU also interested in an MOU, but need real money.
- VLASS-II priorities:
- Validate pipeline for 1" cubes with tapering (SECC)
- Looked into varying cube plane sizes/pixel scales, not much gain.
- Testing tapering to ~2-3".
- Validate and profile hpgawp (GPU) pipeline locally (L4s)
- GPU gridder working in tclean alone, issues with VIP script.
- Waiting for 6.7.1 CASA branch so we can get it into the pipeline.
- GPU gridder working in tclean alone, issues with VIP script.
- Validate pipeline for 1" cubes with tapering (SECC)
- Planning for VLASS-II
- PIPE-2423 - Getting issue details... STATUS
- Get awphpg running on different GPUs (specifically H100 Grace Hoppers at TACC).
- VLASS manager changes/data origin setup.
- K Scott has "read only" version running, need to get products back.
- 30 GPUs arrived, 3 installed in nmpost091-093, can run 27 at once in NMTech.
- Ep3 QL catalog - Dillon to write-up paper with catalogs, we will put them on VLASS website.
- Steve published HIPS to CDS, should appear in e.g. Aladin, probably ESASky too in a few hours.
- Will thus appear in CARTA-5 when registered, should do a joint eNews.
- New Ep1+2 catalog for SECI masks look good, implementation: PIPE-2485 - Getting issue details... STATUS
- Calibrator uv ranges - Tierra testing.
- 2.2 SE calibration begun for EBs that need uv-ranges set
- compressionantexclude - James testing
- Rick Perley checking 3C138 for pol angle changes after it flared.
- Update from Rick: ~2 deg change only
- AOB